OpenStudy (loser66):

How to find 5 consecutive numbers that are not primes. Please, help.

OpenStudy (alexandervonhumboldt2):

wait

OpenStudy (anonymous):

\[5! + 1\]\[5! + 2\]\[\cdots \]\[5! + 5\]

OpenStudy (dayakar):

numbers between 23 and 29

OpenStudy (anonymous):

You can do this in general for any \(n\). If you're given some \(n\), you can try working with the following numbers:\[(n+1)!+2, \ \ (n+1)! + 3, \cdots,\ \ (n+1)! + n+1\]

OpenStudy (ikram002p):

Use the factorial trick as what bea did

OpenStudy (anonymous):

Thus there is always a sequence of \(n\) consecutive nonprime numbers.

OpenStudy (loser66):

@dayakar It is easy to check it out from prime list. but it is not the way to solve @Beauregard if n =1, then (1+1)! = 2! = 2 +2 =4, and 2 +3 =5 is a prime.

OpenStudy (anonymous):

No no, \(n\) is the sequence length. Like in your question \(n = 5\).

OpenStudy (loser66):

got you. Thank you.

OpenStudy (ikram002p):

N can be anything larger or equal 5.

OpenStudy (anonymous):

Oh, you're right. lol

OpenStudy (ikram002p):

=)

OpenStudy (anonymous):

Also the answer I gave initially does not follow the general rule but it is still correct, thankfully.

OpenStudy (anonymous):

The \(n! + 1\) term was a little dangerous as it could have been prime. It is prime for \(n = 1, 2, 3\) so luck saved me.